### v2

@@ 3.1 Strongly Shill-Proof Auctions
Public auctions are common in practice; from open air fish markets, to auctions on eBay, participants often can see every action other bidders take before choosing what to do. [^26] Non-examples of public auctions include the FCC spectrum auctions, where bidders typically only learn information on other bidders’ actions in rounds (see Milgrom and Segal (2017) for more information). Now that we have restricted the possible information structure, we are in a position to present our main result:
@@
A public and optimal auction is strongly shill-proof if and only if it is the Dutch auction with reserve price $\rho^{*}$.
@@
The Dutch auction is strongly shill-proof because any shill bid immediately ends the auction and in that case there would be no transfers from other bidders. To gain an intuition on why uniqueness result holds, we note that the key property of the Dutch auction is that any bid immediately ends the auction. Indeed, for all other auction formats, there exists at least one history such that a bidder can indicate her value is strictly greater than 00 without the auction ending immediately. Such an action has two possible effects: (1) ex-interim increasing the effective reserve price and (2) making it appear as if there is more competition for the item, therefore causing less bid shading. The fact that the auction is public and optimal implies that with a higher reserve price, the transfer from the winner must be higher. [^27] If the auction were not public, then real bidders’ beliefs might not change ex-interim and so their transfers might not change either. For example, in a sealed first-price auction, a shill bidder’s actions have no effect on the transfers of other bidders. We can then conclude that such an auction is not strongly shill-proof because we can consider a valuation vector that generates such a history and have the bidder who can indicate her value is greater than 00 be a shill bidder.
To further understand how the public assumption helps pin down a unique extensive form, please refer to Example O.2 in the Online Appendix for a discussion of single-action, first-price auctions and why they are weakly shill-proof, but not strongly shill-proof.
3.2 Weakly Shill-Proof and Efficient Auctions
To dive further into analyzing when a Dutch auction is needed, we now turn to discussing efficient auctions. While the literature has primarily focused on optimal auctions, efficient auctions are important to consider in many cases. One example is two-sided marketplaces, where the auctioneer/market designer and the seller are different entities and may have different objectives. The designer may be interested in allocating goods efficiently while sellers are trying maximize revenue. Our next result shows that in order for an auction to be weakly shill-proof and efficient robustly to the prior on the value distribution, part of its game tree must be a Dutch auction. In particular, the auction must be a semi-Dutch auction:
@@ 3.2 Weakly Shill-Proof and Efficient Auctions
An efficient auction $(\tilde{x}^{E},\tilde{t},\mu,\xi_{0})$ is a semi-Dutch auction with cutoff $\rho^{*}$ if for any $v$ such that $\max_{i}\left\{v_{i}\right\}<\rho^{*}$:
@@
    $\widecheck{V}=\left\{w:w<\rho^{*}\right\}^{N}$ is reached and
@@
    $\mu(V,\xi)=\mu^{D}(V,\xi)$ for any player $\xi$ and possible values $V\subseteq\widecheck{V}$ where $\mu^{D}$ is the Dutch auction menu rule from Definition 3.1.282828Technically, $\mu^{D}(V,\xi)$ is defined only for optimal auctions; however, it is well-defined for efficient auctions if we redefine $\vec{\xi}_{\mathsf{L}}$ as $\vec{\xi}_{\mathsf{L}}=\begin{cases}\left(\cdot,\tilde{\xi}\right)=\max_{\rhd}\left\{\left(\overline{{V}}_{i},i\right):i\neq\xi\right\}&\exists i\neq\xi\text{ such that }\overline{{V}}_{i}>0\\ \emptyset&\text{otherwise}\end{cases}.$ (1)
[…]
Theorem 3.6.
For every public and efficient auction that is not a semi-Dutch auction with cutoff $\rho^{*}$, there exists a regular value distribution with optimal reserve $\rho^{*}$ under which the auction is not weakly shill-proof.
The key step in the proof of Theorem 3.6 resembles the proof of Theorem 3.4—in any non-Dutch auction, shill bidders can ex-interim “raise the reserve price” by changing their actions. However, given that we are interested in weak shill-proofness instead of strong shill-proofness, we have to examine shill bidders’ incentives when we take expectations over real bidders’ values instead of conditioning directly on their values. Regularity implies that above $\rho^{*}$, shill bidders do not have an incentive to shill bid in auction formats such as the English auction (see Section 1.1). However, below $\rho^{*}$, we can always find a regular distribution such that the ex-interim expected value of raising the reserve price is always positive. In the Appendix, we construct the claimed sub-class of regular distributions (see Definition B.3). Informally, the atoms of the value distribution have to be far enough apart so that raising the reserve price a single “level” generates a large amount of additional revenue. So, below $\rho^{*}$, the auction must resemble the Dutch auction; the class of all such auctions is precisely all semi-Dutch auctions with cutoff $\rho^{*}$. In the Online Appendix (Example O.3), we provide an example of an auction format that is weakly shill-proof for some value distributions but not others; and the following example presents a real-world setting that roughly fits the premises of Theorem 3.6 where a semi-Dutch auction (that is not a Dutch auction) is used:
@@
The Honolulu-Sydney auction plausibly fits the technical assumptions made in Theorem 3.6: The auctions are public, as they take place in person and all bidders can see other bidders’ actions. Market participants are interested in efficient outcomes because the goods are perishable and there are positive disposal costs for the sellers. We do not mean to imply that the Honolulu-Sydney auction was instituted precisely because it is shill-proof, but we highlight it as further evidence that in markets where it is difficult to monitor shill bidding, shill-proof mechanisms often arise.
As a final observation in this section, let us note that if instead of allowing the auction format to treat bids above and below the optimal reserve differently, we instead required the auction format to treat all bids identically, then the only public, efficient, and weakly shill-proof auction is the Dutch auction.
@@
Observe that a semi-Dutch auction with cutoff $\rho^{*}=v^{M}$ is simply a Dutch auction. Then, apply Theorem 3.6 for a regular distribution with optimal reserve $\rho^{*}=v^{M}$. ∎
4 Weakly Shill-Proof and Strategy-Proof Auctions
We have shown that the only optimal auction in which it is an ex-post strategy for shill bidders not to shill (strong shill-proofness) and an equilibrium for real bidders is the Dutch auction. We now investigate the reverse question: what optimal auctions have an ex-post strategy for real bidders (ex-post incentive compatiblility) and an equilibrium for shill bidders not to shill (weak shill-proofness)? Before we explore that question, let us formally define ex-post incentive compatiblility:
Definition 4.1.
An auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$ is ex-post incentive compatible if it is a ex-post strategy for real bidders to report their values truthfully: for all $i$, $v$ and $v_{i}^{\prime}$,
$\tilde{x}(v)v_{i}-\tilde{t}(v)\geq\tilde{x}(v_{i}^{\prime},v_{-i})v_{i}-\tilde{t}(v_{i}^{\prime},v_{-i}).$
@@ 4 Weakly Shill-Proof and Strategy-Proof Auctions
$\tilde{t}^{2}_{i}(v)=\tilde{x}_{i}^{*}(v)\cdot\max\left\{\rho^{*},\text{second-highest value in $\{v_{1},\ldots,v_{N}\}$}\right\}.$
@@
The English auction with reserve price $\rho^{*}$ is defined as the auction with the optimal allocation rule $\tilde{x}^{*}$, second-price transfer rule $\tilde{t}^{2}$, initial player $\left(\cdot,\xi_{0}\right)=\min_{\rhd}\left\{\left(0,i\right)\right\}_{i\in B}$, and menu
$\mu^{E}(V,\xi)$ =$=\left\{\left(W_{\mathsf{L}},\vec{\xi}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\vec{\xi}_{\mathsf{H}}\right)\right\},$
where $W_{\mathsf{L}}=\left\{v\in V_{\xi}:v<\rho^{*}\right\}\cup\left\{\underline{{V}}_{\xi}\right\},W_{\mathsf{H}}=V_{\xi}\backslash W_{\mathsf{L}},$
$\vec{\xi}_{\mathsf{L}}=\vec{\xi}_{\mathsf{H}}=\begin{cases}\left(\cdot,\tilde{\xi}\right)=\min_{\rhd}\left\{\left(\underline{{V}}_{i},i\right):i\neq\xi,\overline{{V}}_{i}=v^{M},\left|V_{i}\right|>1\right\}&\overline{{V}}_{-\xi}=v^{M}\\ \emptyset&\text{otherwise}\end{cases}.$
@@
The English auction with reserve price $\rho^{*}$ is weakly shill-proof, ex-post incentive compatible, and optimal. Depending on the information bidders have when taking actions, it can also easily be made strategy-proof. See Section 6.2 for a discussion of how our results hold in dominant strategies instead of ex-post strategies.
The English auction is not the only ex-post incentive compatible and weakly shill-proof auction. While the English auction is used frequently, one drawback is that it is “slow”—each bidder can be queried on their willingness-to-pay on the order of $M$ times. Specifically, let $Q^{E}(F)=\left|\left\{k:v^{k}\geq\rho^{*}\right\}\right|-1$ be the worst-case number of times a bidder must be queried. To explore if there are weakly shill-proof and ex-post incentive compatible auctions that require fewer rounds of communication, we introduce a natural “compression” of the English auction. The ascending, screening auction, comprises the following two phases:
@@
    An English auction is run from $\rho^{*}$ to some $v^{Y}$.
@@
    If necessary, a second-price auction is then run among players who have not dropped out before the value level of $v^{Y}$.
@@
The ascending, screening auction with screen level $v^{Y}$ is defined by the optimal allocation rule $\tilde{x}^{*}$, second-price transfer rule $\tilde{t}^{2}$, initial player $\left(\cdot,\xi_{0}\right)=\min_{\rhd}\left\{\left(\overline{{V}}_{i},i\right)\right\}$, and menu
$\mu(V,\xi)=\begin{cases}\mu^{E}(V,\xi)&\exists i\text{ such that }\underline{{V}}_{i}<v^{Y}\text{ and }\overline{{V}}_{i}=v^{M}\\ \left\{\left(\left\{v^{k}\right\},\vec{\xi}\ \right)\right\}_{k\in\left\{Y,Y+1,\ldots,M\right\}}&\text{otherwise}\end{cases},$
where (,$\text{where }\left(\cdot,\vec{\xi}\ \right)=\max_{\rhd}\left\{(0,i):\left|V_{i}\right|>1,\underline{{V}}_{i}=v^{Y}\right\}\text{ and }\vec{\xi}^{\,M}=\emptyset.$
This auction reduces the maximum number of times each bidder can be queried to $Q^{AS,Y}(F)=\left|\left\{k:\rho^{*}\leq v^{k}\leq v^{Y}\right\}\right|+1.$ Because the transfer rule is $\tilde{t}^{2}$, the ascending, screening auction is ex-post incentive compatible and optimal.

### v3

@@ 3.3 Dutch Auctions
Public auctions are common in practice: from open air fish markets, to auctions on eBay, participants often can see every action other bidders take before choosing what to do. [^22] Non-examples of public auctions include the FCC spectrum auctions, where bidders typically only learn information on other bidders’ actions in rounds (see Milgrom and Segal (2017) for more information). A second interpretation of an auction being public is that all information is totally leakable: If an auction is strongly shill-proof and public, then shill bidders can credibly signal to all other bidders any actions they have taken in the auction and it will not affect the outcome. In general, we believe that it is natural to assume that shill bidders actions are visible in settings where considering shill bidding is important. Otherwise, shill bidding can only manipulate the outcome ex-post, which differs from the normal interpretation of shill bidding and its applications.
Next, we define the Dutch auction with reserve price $\theta^{\rho}$. The Dutch auction is defined as the auction which begins by offering each bidder $i$ the item at $b^{1}_{i}(\theta^{M})$, and then if no bidder chooses to buy the item at that price, the item is offered for $b^{1}_{i}(\theta^{M-1})$ and so on until either a bidder has chosen to buy the item or the price to be offered drops below $b^{1}_{i}(\theta^{\rho})$. Note that the optimal Dutch auction is the Dutch auction with reserve price $\theta^{\rho^{*}}$. We consider only orderly auctions and therefore, at each price level, bidders are offered the opportunity to buy the item in priority order. Formally:
Definition 3.5.
The Dutch auction with reserve price $\theta^{\rho}$ is defined by the allocation rule $\tilde{x}^{\rho}$, first-price transfer rule $\tilde{t}^{1}=\tilde{x}^{\rho}\cdot b^{1}$, initial player $\left(\cdot,\xi_{0}\right)=\max_{\rhd}\left\{\left(0,i\right)\right\}$, and menu
$\mu^{D}_{\rho}(\Theta,\xi)$ $=\left\{\left(W_{\mathsf{L}},\vec{\xi}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\vec{\xi}_{\mathsf{H}}\right)\right\},$
where $W_{\mathsf{H}}=\left\{\overline{{\Theta}}_{\xi}\right\},W_{\mathsf{L}}=\Theta_{\xi}\backslash\left\{\overline{{\Theta}}_{\xi}\right\},\vec{\xi}_{\mathsf{H}}=\emptyset,\text{ and }$
$\vec{\xi}_{\mathsf{L}}=\begin{cases}\left(\cdot,\tilde{\xi}\right)=\max_{\rhd}\left\{\left(\overline{{\Theta}}_{i},i\right):i\neq\xi\right\}&\exists i\neq\xi\text{ such that }\left|\Theta_{i}\right|>1\text{ and }\overline{{\Theta}}_{i}\geq\theta^{\rho}\\ \emptyset&\text{otherwise}\end{cases}.$
@@
A Dutch auction with any reserve price is strong shill-proof. Furthermore, if a public, orderly and optimal auction is strongly shill-proof, then it is the Dutch auction with reserve price $\theta^{\rho^{*}}$.
Proof Sketch..
@@
The Dutch auction is strongly shill-proof because any shill bid immediately ends the auction—and in that case there would be no transfers from other bidders. To gain an intuition for why uniqueness holds, note that the key defining property of the Dutch auction is that any bid immediately ends the auction. Then, we prove the result in four cases. First, we consider an auction where the player rotation differs from a Dutch auction; we then show that if a shill bidder indicates she has the highest possible remaining value, this has the effect of ex-interim increasing the effective reserve price. Since the auction is public and only the highest value is allocated, when the types are realized so that some bidder with a higher priority order has the same value (and non-shill bidders have a low value), raising the reserve must increase revenue from the winning bidder. [^23] If the auction were not public, then real bidders’ beliefs might not change ex-interim and so their transfers might not change either. For example, in a sealed first-price auction, a shill bidder’s actions have no effect on the transfers of other bidders. The next case considers what happens when bidders are queried in priority order, and the minimum of a partition is greater than the reserve price, but less than the highest value of other bidders. This raises the ex-interim reserve price and so by the same argument as above, shill bidding is profitable. In the third case, we consider what happens when the minimum in the partition is equal to the highest value of other bidders. By construction, this can only occur when a bidder is the lowest priority and so, applying our orderliness assumption, this raises the ex-interim reserve to be the highest type still possible in equilibrium for other bidders without immediately allocating the good to the shill bidder. We can then apply the same argument as the first two cases to conclude that shill bidding would be profitable in this case. In the final case, we consider an auction where the minimum in the partition is below the reserve price. In this case, the ex-interim reserve has not changed, so the logic from the previous cases does not apply. Instead, shill bidding in this case makes it appear as if there is more competition for the item, and we prove that this causes “less bid shading,” i.e., higher final transfers from the winner. ∎
Observe that although the standard intuition about shill bidding is that its’ purpose is to influence other bidders’ perception about the common value for an item, Theorem 3.6 shows that just being able to influence other bidders perception about the probability they win the item without directly winning the item in a single case is enough to limit the number of possible auction formats to just Dutch auctions. Moreover, under public information, such a case always exists for non-Dutch auctions. The public information structure is the most informative, while a static information structure is the least informative. We show in the Supplemental Appendix that as the information structure becomes less informative, the number of strongly shill-proof auctions weakly increases (Proposition E.3).
4 Weakly Shill-Proof Auctions
In Lemma 3.1, we demonstrate that strong shill-proofness uniquely pins down the transfer rule; and in Theorem 3.6, we demonstrate that once we consider public auctions, we uniquely pin down the extensive form as well. In this section, we turn to studying our more permissive definition, weak shill-proofness, and first show the limits of possible extensive forms that are weakly shill-proof and then exhibit extensive-form auctions that are weakly shill-proof and strategy-proof, including a new auction format that can be arbitrarily faster than the English auction.
4.1 Trilemma
In this subsection, we prove that it is impossible to find an auction that is weakly shill-proof, strategically simple for even one of the real bidders, and fast in the sense that each bidder takes exactly one action. To make this trilemma as general as possible, we do not even require the auction to be optimal or efficient. Instead, we are only using condition (iii) of our orderliness definition (Definition 3.2), which requires that a bidder who does not have the highest type must not win the item. Note that such a condition is without loss for any symmetric and deterministic mechanism.
We begin by defining a single-action auction. An auction is considered single-action when each bidder takes precisely one action in the auction (under all possible histories). More formally, for any $h_{N}\in Z$, let $h_{\emptyset}\prec h_{1}\prec\ldots h_{N-1}\prec h_{N}$ be the sequence of preceding histories. Then, an auction is single-action if for all $i\in B$, there exists a unique $n\leq N$ such that $i=P(h_{n})$. Without loss, we label the bidders $1,\ldots,N$, in the order that they move and label the action taken by bidder $i$ as $a_{i}$. [^24] The bidder ordering and therefore the labeling can be endogenous to actions taken.
For expositional purposes, instead of tracking the information set $\mathcal{I}_{i}$ of a bidder $i\in R$, we assume that the information $i$ has when taking an action is a signal $s_{i}\in\mathcal{S}_{i}$. This signal is generated via a deterministic function $\psi_{i}:\left(\bigtimes_{j<i}A_{j}\right)\to\mathcal{S}_{i}$ called an experiment. [^25] Abusing notation, we also sometimes take $\psi_{i}:\vartheta^{i-1}\to\mathcal{S}_{i}$, i.e., the experiment maps values to signals instead of actions. For notational convenience, we assume that $\psi_{i}$ is surjective for all $i$. We can think of the experiment as a garbling of the previous bidders’ actions—the experiment can pool together multiple actions from previous bidders to a single signal and so a signal is not always perfectly informative of previous actions. We can recover the public setting with a fully informative experiment, i.e., $\psi=\mathrm{Id}$, the identity mapping. We can capture classical static game settings via an uninformative experiment that always returns the same output, $\psi=\emptyset$. We use $\psi^{-1}_{i}(s_{i})$ to denote the set of $\theta_{-i}$ that are possible from the perspective of bidder $i$ given its signal.
A revelation principle holds in this setting: for any single-action auction, we can define the direct allocation and transfer rules as $\tilde{x}(\theta)$ and $\tilde{t}(\theta)$, respectively, with the appropriate incentive compatibility and individual rationality constraints for real bidders (Lemma C.1 in the Appendix), and appropriate IC constraints for weak and strong shill-proofness (Lemmatta C.2 and E.13, respectively, in the Appendix).
To finish defining all the terms necessary for our main result of this section, we formally define ex-post incentive compatibility and then give a weaker notion of ex-post incentive compatiblility in the single-action auction setting: ex-post incentive compatiblility for at least a single bidder.
@@ 4.1 Trilemma
An auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$ is ex-post incentive compatible if it is an ex-post strategy for real bidders to report their values truthfully: for all real bidders $i\in R$, $\theta$, and $\theta_{i}^{\prime}$,
$\tilde{x}(\theta)\cdot\theta_{i}-\tilde{t}(\theta)\geq\tilde{x}(\theta_{i}^{\prime},\theta_{-i})\cdot\theta_{i}-\tilde{t}(\theta_{i}^{\prime},\theta_{-i}).$
Definition 4.2.
A single-action auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$ is mildly ex-post incentive compatible if there exists a real bidder $i<N$ such that truthfulness is an ex-post strategy for $i$ conditional on the realization of her signal: there exists bidder $i<N$, such that for all $\theta_{i}$, $\theta_{i}^{\prime}$, $s_{i}$, and $\theta_{-i},\theta_{-i}^{\prime}\in\psi_{i}^{-1}(s_{i})$: $\tilde{x}_{i}(\theta)\cdot\theta_{i}-\tilde{t}_{i}(\theta)\geq\tilde{x}_{i}\left(\theta^{\prime}\right)\cdot\theta_{i}-\tilde{t}_{i}\left(\theta^{\prime}\right)$. [^26] We exclude the last bidder who takes an action from our definition because a take-it-or-leave-it offer to that bidder can be optimal and ex-post incentive compatible. Theorem 4.3 would still hold if we instead defined mild ex-post incentive compatiblility to mean ex-post incentive compatible for at least two bidders.
Theorem 4.3.
There exists no orderly auction that is single-action, mildly ex-post incentive compatible, and weakly shill-proof.
Proof Sketch.
Consider any real bidder $i<N$. By weak shill-proofness, the transfer from bidder $i$, conditional on winning (or losing) the auction, is invariant to the values of bidders who take actions after her (Lemma C.3 in the Appendix). If this were not the case, then if every bidder $j>i$ is a shill bidder, the shill bidders would report the values that would maximize the transfer from the winning bidder. By mild ex-post incentive compatiblility, the transfer from bidder $i$, conditional on winning the auction, is invariant to her value (Lemma C.4 in the Appendix). This is because if there were multiple winning reports with different transfer amounts, only the smallest transfer amount would make truthful reporting of the value an ex-post strategy. So, in every single-action, optimal auction, the transfer from the winning bidder $i$, can depend only on the values reported by bidders before $i$. But, this means that if a bidder has positive utility for winning the item (as would be the case if $\theta_{i}>\theta_{j}$ for all $j<i$), then she should report $\theta^{M}$ to maximize the probability of winning (without changing the transfer paid upon winning). Thus, the auction must treat bidder $i$ as if she reported $\theta^{M}$, which violates the allocation rule of an optimal auction. ∎
Observe that while there are no auctions at the intersection of all three conditions in Theorem 4.3, we can find multiple auctions at the intersection of any two of those conditions. For example, assuming an IPV and regular (Definition 4.4) environment, the second-price auction is single-action and mildly ex-post incentive compatible; the first-price auction is single-action and weakly shill-proof; and the English auction is mildly ex-post incentive compatible and weakly shill-proof. None of the examples just described are unique; there exist other single-action and mildly ex-post incentive compatible auctions, single-action and weakly shill-proof auctions, and mildly ex-post incentive compatible and weakly shill-proof auctions.
4.2 Other Weakly Shill-Proof Auctions
Now that we have shown an impossibility result regarding weakly shill-proof auctions, in this subsection, we will provide some characterization results for weakly shill-proof auctions. Recall from the introduction that even in an IPV environment, an English auction need not be weakly shill-proof when the type distribution is irregular. So, for tractability, for the rest of this subsection, we suppose that bidders have IPV types and regular type distributions. We use the definition of regularity for discrete types found in Elkind (2007):
Definition 4.4.
A distribution $F$ is regular if for all $k$, the virtual value $\varphi^{k}=\theta^{k}-(\theta^{k+1}-\theta^{k})\frac{1-F(\theta^{k})}{f(\theta^{k})}$ is non-decreasing.
4.2.1 Weakly Shill-Proof and Efficient Auctions
To further analyze when concerns about shill-bidding lead to Dutch auctions, we now examine the case of efficient auctions. Shill bidding is important to consider in the efficient auction context because the designer may be interested in allocating goods efficiently even while sellers are trying maximize revenue. For example, in two-sided marketplaces like eBay, the auctioneer/market designer places positive welfare weight on both buyers and sellers, whereas sellers typically do not—and sellers may certainly try to shill bid in these settings, as discussed in the Introduction. Our next result shows that in order for an auction to be robustly weakly shill-proof and efficient, part of its game tree must be a Dutch auction. If an auction is not robustly weakly shill-proof, we mean that we can find a value distribution such that the auction is not weakly shill-proof. More formally, if the auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$ is parameterized by the optimal reserve $\theta^{\rho^{*}}$, the number of atoms below the reserve $\underline{M}$ and the number of atoms weakly above the reserve $\overline{M}$, [^27] Here, $\underline{M}+\overline{M}=M$. then the following result holds.
Definition 4.5.
An efficient auction $(\tilde{x}^{E},\tilde{t},\mu,\xi_{0})$ is a semi-Dutch auction with cutoff $\theta^{\rho^{*}}$ if for any $v$ such that $\max_{i}\left\{v_{i}\right\}<\theta^{\rho^{*}}$:
@@ 4.2.1 Weakly Shill-Proof and Efficient Auctions
    $\widecheck{\Theta}=\left\{w:w<\theta^{\rho^{*}}\right\}^{N}$ is reached; and
@@
    $\mu(\Theta,\xi)=\mu^{D}(\Theta,\xi)$ for any player $\xi$ and possible values $\Theta\subseteq\widecheck{\Theta}$ where $\mu^{D}$ is the menu rule for the Dutch auction with reserve price 0 from Definition 3.5.
Proposition 4.6.
For every public and efficient auction that is not a semi-Dutch auction with cutoff $\theta^{\rho^{*}}$, there exists a regular value distribution with optimal reserve $\theta^{\rho^{*}}$ under which the auction is not weakly shill-proof.
The key step in the proof of Proposition 4.6 resembles the proof of Theorem 3.6—in any non-Dutch auction, shill bidders can ex-interim “raise the reserve price” by changing their actions. However, given that we are interested in weak shill-proofness instead of strong shill-proofness, we have to examine shill bidders’ incentives when we take expectations over real bidders’ values instead of conditioning directly on their values. Regularity implies that above $\theta^{\rho^{*}}$, shill bidders do not have an incentive to shill bid in auction formats such as the English auction (see Section 1.1). However, below $\theta^{\rho^{*}}$, we can always find a regular distribution such that the ex-interim expected value of raising the reserve price is always positive; in particular, we can find a value distribution where the atoms are far enough apart that raising the reserve price a single “level” generates a large amount of additional revenue. In the Appendix, we construct the claimed sub-class of regular distributions (see Definition C.5). So, below $\theta^{\rho^{*}}$, the auction must resemble the Dutch auction; the class of all such auctions is precisely all semi-Dutch auctions with cutoff $\theta^{\rho^{*}}$. In the Supplemental Appendix (Example E.5), we provide an example of an auction format that is not semi-Dutch and that is weakly shill-proof for some value distributions but not others; and the following example presents a real-world setting that roughly fits the premises of Proposition 4.6 where a semi-Dutch auction (that is not a Dutch auction) is used.
@@
The Honolulu-Sydney auction plausibly fits the technical assumptions made in Proposition 4.6: The auctions are public, as they take place in person and all bidders can see other bidders’ actions. Market participants are interested in efficient outcomes because the goods are perishable and there are positive disposal costs for the sellers. We do not mean to imply that the Honolulu-Sydney auction was instituted precisely because it is shill-proof, but we highlight it as further evidence that in markets where it is difficult to monitor shill bidding, shill-proof mechanisms may arise.
Note that Proposition 4.6 relies on the auction format being able to condition on the true, optimal reserve. If we instead require the auction format to be completely prior-independent, then the only public, efficient, and weakly shill-proof auction is the Dutch auction.
@@
Observe that a semi-Dutch auction with cutoff $\theta^{\rho^{*}}=\theta^{M}$ is simply a Dutch auction. Then, apply Proposition 4.6 for a regular distribution with optimal reserve $\theta^{\rho^{*}}=\theta^{M}$. ∎
4.2.2 Weakly Shill-Proof and Strategy-Proof Auctions
We have shown that the only optimal auction with a feasible equilibrium for real bids and an ex-post strategy for shill bidders not to shill (strong shill-proofness) is the Dutch auction. We now investigate the dual question: what optimal auctions have an ex-post strategy for real bidders (ex-post incentive compatiblility) and an equilibrium under which no shill bidding occurs (weak shill-proofness)?
@@ 4.2.2 Weakly Shill-Proof and Strategy-Proof Auctions
$\tilde{t}^{2}_{i}(\theta)=\tilde{x}_{i}^{*}(\theta)\cdot\max\left\{\theta^{\rho^{*}},\text{second-highest value in $\{\theta_{1},\ldots,\theta_{N}\}$}\right\}.$
@@
The English auction with reserve price $\theta^{\rho^{*}}$ is defined as the auction with the optimal allocation rule $\tilde{x}^{*}$, second-price transfer rule $\tilde{t}^{2}$, initial player $\left(\cdot,\xi_{0}\right)=\min_{\rhd}\left\{\left(0,i\right)\right\}_{i\in B}$, and menu
$\mu^{E}(\Theta,\xi)$ $=\left\{\left(W_{\mathsf{L}},\vec{\xi}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\vec{\xi}_{\mathsf{H}}\right)\right\},$
where $W_{\mathsf{L}}=\left\{\theta\in\Theta_{\xi}:\theta<\theta^{\rho^{*}}\right\}\cup\left\{\underline{{\Theta}}_{\xi}\right\},W_{\mathsf{H}}=\Theta_{\xi}\backslash W_{\mathsf{L}},$
$\vec{\xi}_{\mathsf{L}}=\vec{\xi}_{\mathsf{H}}=\begin{cases}\left(\cdot,\tilde{\xi}\right)=\min_{\rhd}\left\{\left(\underline{{\Theta}}_{i},i\right):i\neq\xi,\overline{{\Theta}}_{i}=\theta^{M},\left|\Theta_{i}\right|>1\right\}&\overline{{\Theta}}_{-\xi}=\theta^{M}\\ \emptyset&\text{otherwise}\end{cases}.$
@@
The English auction with reserve price $\theta^{\rho^{*}}$ is weakly shill-proof, ex-post incentive compatible, and optimal (when the value distribution is regular). Depending on the information bidders have when taking actions, that auction can also be made dominant-strategy. [^29] See Section 6.2 for a discussion of how our results translate into dominant-strategy contexts.
The English auction is not the only ex-post incentive compatible and weakly shill-proof auction. While the English auction is used frequently, one drawback is that it is “slow”—each bidder can be queried on their willingness-to-pay on the order of $M$ times. Specifically, let $Q^{E}(F)=M-\rho^{*}+1$ be the worst-case number of times a bidder must be queried. To explore whether there are weakly shill-proof and ex-post incentive compatible auctions that require fewer rounds of communication, we introduce a natural “compression” of the English auction that comprises the following two phases:
@@
    An English auction is run from $\theta^{\rho^{*}}$ to some $\theta^{Y}$.
@@
    If necessary, a second-price auction is then run among players who have not dropped out before the value level of $\theta^{Y}$.
@@
The ascending, screening auction with screen level $\theta^{Y}$ is defined by the optimal allocation rule $\tilde{x}^{*}$, second-price transfer rule $\tilde{t}^{2}$, initial player $\left(\cdot,\xi_{0}\right)=\min_{\rhd}\left\{\left(\overline{{\Theta}}_{i},i\right)\right\}$, and menu
$\mu(\Theta,\xi)=\begin{cases}\mu^{E}(\Theta,\xi)&\exists i\text{ such that }\underline{{\Theta}}_{i}\leq\theta^{Y}\text{ and }\overline{{\Theta}}_{i}=\theta^{M}\\ \left\{\left(\left\{\theta^{k}\right\},\vec{\xi}^{\,k}\right)\right\}_{k\in\left\{Y+1,\ldots,M\right\}}&\text{otherwise}\end{cases},$
$\text{where }\left(\cdot,\vec{\xi}^{\,k}\right)=\max_{\rhd}\left\{(0,i):\left|\Theta_{i}\right|>1,\underline{{\Theta}}_{i}=\theta^{Y}\right\}\text{ for $k<M$ and }\vec{\xi}^{\,M}=\emptyset.$
The ascending, screening auction reduces the maximum number of times each bidder can be queried to $Q^{AS,Y}(F)=Y-\rho^{*}+2.$ Because the transfer rule is $\tilde{t}^{2}$, the ascending, screening auction is ex-post incentive compatible and optimal.